OpenStudy (anonymous):

Iam completely lost on doing the problem can someone plz explain to me how to do it. simplify the given expression sqrt-64/(7-6i)-(2-2i)

terenzreignz (terenzreignz):

Sure :) \[\Large \frac{\sqrt{-64}}{(7-6\color{red}i)-(2-2\color{red}i)}\]

terenzreignz (terenzreignz):

So, what is the square root of 64?

OpenStudy (anonymous):

i know i have to distribute

terenzreignz (terenzreignz):

distribute? Like in the denominator? Well, go ahead, then :)

OpenStudy (anonymous):

but he sqrt of 64 is 8

terenzreignz (terenzreignz):

That's right... but it's -64 What's the square root of -64?

OpenStudy (anonymous):

-8?

terenzreignz (terenzreignz):

No... because -8 times -8 is still 64 :) We're talking about complex numbers anyway, the square root of a negative number is i times the square root of the number. for instance \[\Large \sqrt{-25}= 5\color{red}i\]

terenzreignz (terenzreignz):

That said, what's the square root of -64?

OpenStudy (anonymous):

\[\sqrt{-64}=8i\]

terenzreignz (terenzreignz):

You catch on quickly :) So let's replace the numerator... \[\Large \frac{8\color{red}i}{(7-6\color{red}i)-(2-2\color{red}i)}\] Now simplify the denominator... you can do that by combining like terms, much like in algebra :)

OpenStudy (anonymous):

8i/14-14i-12i+12i^2

terenzreignz (terenzreignz):

OH... there is no multiplication involved (yet) It's just adding and subtracting

terenzreignz (terenzreignz):

Distribute the negative sign first, to make things easier. \[\Large \frac{8\color{red}i}{(7-6\color{red}i)\color{green}-(2-2\color{red}i)}\]

OpenStudy (anonymous):

-2+2i

terenzreignz (terenzreignz):

Great :) so it now becomes \[\Large \frac{8\color{red}i}{7-6\color{red}i-2+2\color{red}i}\] combine like terms in the denominator, you get?

OpenStudy (anonymous):

yes 5-4i

terenzreignz (terenzreignz):

Brilliant :) \[\Large \frac{8\color{red}i}{5-4\color{red}i}\] Now, much like radicals, people don't like imaginary units in the denominator...

OpenStudy (anonymous):

so i multipliy by i correct?

terenzreignz (terenzreignz):

A way to get rid of that imaginary unit is multiply both the numerator and denominator by the CONJUGATE of the denominator. You know what a conjugate is? :)

OpenStudy (anonymous):

yes means you use the opposite sign

OpenStudy (anonymous):

(5+4i)

terenzreignz (terenzreignz):

Very nice :)

terenzreignz (terenzreignz):

Multiply both the numerator and denominator by that conjuate \[\Large \frac{8\color{red}i\cdot(5+4\color{red}i)}{(5-4\color{red}i)(5+4\color{red}i)}\]

terenzreignz (terenzreignz):

conjugate*

OpenStudy (anonymous):

40i+32i^2/25i

terenzreignz (terenzreignz):

Whoa now... work on that denominator again... \[\Large (a-b)(a+b)=a^2-b^2\]

OpenStudy (anonymous):

ok wil it be -3

terenzreignz (terenzreignz):

Nope :) Look at it, it's a sum and difference... Use this form \[\Large (a-b)(a+b)=a^2-b^2\] where a = 5 b = 4i

OpenStudy (anonymous):

(5-4i)(5+4i)=25-16i^2

terenzreignz (terenzreignz):

Good, now just remember that \[\Large \color{red}i^2 = -1\]

OpenStudy (anonymous):

ok 32+40i/-1

terenzreignz (terenzreignz):

What? LOL don't be so hasty. \[\Large 25 - 16\color{red}i^2 = 25 - 16(\color{red}{-1})\] simplify

OpenStudy (anonymous):

i am confused:/

OpenStudy (anonymous):

will it be25-16i=25+16

terenzreignz (terenzreignz):

Precisely :) Remember, the whole reason we multiplied both the numerator and denominator by the conjugate...

terenzreignz (terenzreignz):

WAS TO GET RID OF THE \(\large \color{red}i\) in the denominator !!!

OpenStudy (anonymous):

oh ok

terenzreignz (terenzreignz):

So you understand it now :) At the numerator, what did you get?

OpenStudy (anonymous):

i do,but the only answer choice i have that is close to this is 32+40i/41

OpenStudy (anonymous):

is that right

terenzreignz (terenzreignz):

Check the numerator...

terenzreignz (terenzreignz):

\[\Large 8\color{red}i(5+4\color{red}i)\] distribute?

OpenStudy (anonymous):

40i+32i^2

OpenStudy (anonymous):

wich gives me -31

terenzreignz (terenzreignz):

Interesting. And it's -32 btw :)

terenzreignz (terenzreignz):

What are the answer choices?

OpenStudy (anonymous):

on sec

OpenStudy (anonymous):

ok. a.-64+40i/-3 b.34+40i/41 c.64+40i/-3 d.-32+40i/41

terenzreignz (terenzreignz):

smh... there's a -32 in choice (d) :P

OpenStudy (anonymous):

thanks for your help i am not very good in math

terenzreignz (terenzreignz):

You actually are. Just need to avoid that spark of carelessness every now and then :)

OpenStudy (anonymous):

lol thanks can you check two more for me to see if i did them correctly? if you have time.

terenzreignz (terenzreignz):

Hit me

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

sqrt-6(2+sqrt8) a.2sqrt6+4sqrt3 b.-2sqrt6+4sqrt3 c.-4sqrt3+2isqrt6 d.4sqrt3+2isqrt6 i chose d

terenzreignz (terenzreignz):

Correct :)

OpenStudy (anonymous):

ok last one

OpenStudy (anonymous):

sqrt-10*sqrt-8 a.-4sqrt5 b.4sqrt5 c.4isqrt5 d.-4isqrt5 i chose c

terenzreignz (terenzreignz):

wrong :) remember that multiplying through radicals is only valid if the radicands are positive, which they aren't :)

OpenStudy (anonymous):

so the 4 has to be a negative correct

terenzreignz (terenzreignz):

Yup :)

OpenStudy (anonymous):

ok thanks so much i feel beteoing these now :)

terenzreignz (terenzreignz):

wait. what's your answer?

OpenStudy (anonymous):

sorry d is my answer

terenzreignz (terenzreignz):

still no :) Look, I'm tired, so I'll run it through you in as much detail as possible. Try to study the steps thoroughly, okay? :)

OpenStudy (anonymous):

ok thank u

terenzreignz (terenzreignz):

\[\Large \sqrt{-10}\sqrt{-8}\]

terenzreignz (terenzreignz):

So, we take out the i's... \[\Large (\color{red}{i}\sqrt{10})(\color{red}i\sqrt{8})\]\[\Large \color{red}i^2\sqrt{10}\sqrt8\] \[\Large \color{red}i^2\sqrt{40}\]

terenzreignz (terenzreignz):

Again, let's not forget that \[\Large \color{red}i^2 = -1\]

terenzreignz (terenzreignz):

Whoops, typo. Sorry.. \[\Large \color{red}i^2\sqrt{\color{blue}{80}}\]

terenzreignz (terenzreignz):

Now... \(\large \color{red}i^2 = -1\) \[\Large (-1)\sqrt{16\cdot 5}\] \[\Large (-1)4\sqrt5\] \[\Huge\color{blue}{-4\sqrt5}\]

OpenStudy (anonymous):

oh ok i see now thanks for ur help

terenzreignz (terenzreignz):

Practice okay? Promise... ;)

OpenStudy (anonymous):

i promise

terenzreignz (terenzreignz):

That's all I need to know :) ---------------------------------- Terence out